How do I write an equation in slope intercept form for the line that passes through (5,6) and is parallel to the line described by 4x - 5y = 20
Answer by nerdybill(7384)

How do I write an equation in slope intercept form for the line that passes through (5,6) and is parallel to the line described by 4x - 5y = 20
.
First, determine slope of
4x - 5y = 20
.
do this by putting it into slope-intercept form y=mx+b:
4x - 5y = 20
-5y = -4x + 20
5y = 4x - 20
y = (4/5)x - 16
slope is 4/5
since the new line is parallel it must have the same slope.
.
Plug the slope (4/5) and the given point (5,6) into the point-slope form:
y - y1 = m(x - x1)
y - 6 = (4/5)(x - 5)
y - 6 = (4/5)x - 4
y = (4/5)x - 4 + 6
y = (4/5)x + 2 (this is what they're looking for)